AshyKashy:

Hello Ya'll, Please Help Me! ;-; c) Use the formula to work out the following pressure problem. i) A stone statue has a weight of 3000N and its base has an area of 2m2 -What is the pressure exerted by the statue on the floor ? ........................................................................................................ ..................................................................................................................... .......................................................................................................... Please, please help, thank you all!♡´・ᴗ・`♡

jimthompson5910:

\(\large \text{Pressure} = \frac{\text{Force}}{\text{Area}}\) For example, if you have 5000 Newtons of force applied to an area of 10 square meters, then the pressure will be: \(\large \text{Pressure} = \frac{5000 \text{ N}}{10\text{ m}^2}\) \(\large \text{Pressure} = \frac{5000 }{10}\frac{\text{ N}}{\text{ m}^2}\) \(\large \text{Pressure} = 500 \text{ Pascals}\) This is just an example and not the answer to your question. Follow these steps to get your answer. Side notes: The unit "Pascals" abbreviates to "Pa" 1 Pascal is equal to 1 Newton of force per square meter.

AshyKashy:

Thank you!